OpenStudy (anonymous):

Where does the graph of y = 3x2 + 14x − 5 cross the x-axis?

OpenStudy (anonymous):

FLVS?

OpenStudy (anonymous):

yes

OpenStudy (campbell_st):

you need to let y = 0.... and then solve the equation 0 = 3x^2 + 14x - 5 this can be factorised

OpenStudy (anonymous):

or you can do the same with your previous problem (wit the firecracker) with the quadratic formula...

OpenStudy (anonymous):

Okay thanks guys

OpenStudy (campbell_st):

an easy method for your question is 3 x - 5 = -15 find the factors of -15 that add to 14...... 15, -1 so then its  take out common factors  which leaves \[(x+5)(3x -1) = 0\] just solve x + 5 = 0 3x -1 = 0 for x